OpenStudy (anonymous):

4/v =3/5 Round your answer to the nearest ten..

OpenStudy (bibby):

Use opposite operations. Multiply both sides by \[\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

I still don't get it @bibby :(

OpenStudy (bibby):

Have you learned about cross multiplication?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

So why can't you approach this problem?

OpenStudy (anonymous):

I'm just confused which numbers am I supposed to multiply them by? How am I supposed to multiply the v by 3/5? @bibby

OpenStudy (bibby):

You don't multiply the v side by 3/5. If you want to just do it without understanding the math, you can multiply across the diagonals. This is what the actual math looks like:\[\frac{ 5 }{3 }*\frac{ 4 }{v } = \frac{ 3 }{ 5 }*\frac{ 5 }{3 }\]

OpenStudy (bibby):

if you can see, when you multiply by the flipped fraction, the right hand side cancels out

OpenStudy (anonymous):

oh.... thank you :) I understand sorta... you have to use the reciprocal? @bibby

OpenStudy (bibby):

Kinda. You use the reciprocal of the side without the variable. You also have to multiply b oth sides by the same number that way it's still correct according to the rules of algebra

OpenStudy (anonymous):

Thanks :) @bibby